# Leopard[Uncaptioned image]: A Vision Language Model for Text-Rich Multi-Image Tasks

### v2

@@ A.2 Prompts
A.3 Details of Table Rendering
To convert the textual table dataset into a multimodal dataset, the JSON or DataFrame format data is transformed into tabular images using Python. We utilize three Python packages, i.e. dataframe_image [^5] https://github.com/dexplo/dataframe_image., pandas [^6] https://pandas.pydata.org/., and matplotlib [^7] https://matplotlib.org/. with various styling to enhance the diversity of the rendered images. To ensure the clarity and legibility of the plotted images, the original data is filtered by excluding any tables that contain more than 20 rows. This threshold was set to maintain the recognizability of the resulting images.
A.4 Qualitative Results

### v3

@@ A.3 Prompts
A.4 Details of Table Rendering
To convert the textual table dataset into a multimodal dataset, the JSON or DataFrame format data is transformed into tabular images using Python. We utilize three Python packages, i.e., dataframe_image [^4] https://github.com/dexplo/dataframe_image., pandas [^5] https://pandas.pydata.org/., and matplotlib [^6] https://matplotlib.org/. with various styling to enhance the diversity of the rendered images. To ensure the clarity and legibility of the plotted images, the original data is filtered by excluding any tables that contain more than 20 rows. This threshold was set to maintain the recognizability of the resulting images.
A.5 Qualitative Results